Acetate, Inc., has equity with a market value of $24 million and debt with a market value of $9.6 million. Treasury bills that mature in one year yield 6 percent per year, and the expected return on the market portfolio is 11 percent. The beta of the company’s equity is 1.25. The company pays no taxes. a.What is the company's debt–equity ratio? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Debt–equity ratio b. What is the company’s weighted average cost of capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Weighted average cost of capital % c. What is the cost of capital for an otherwise identical all-equity company? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Cost of capital %
Answer #1

a) D/E ratio = 9.6 / 24 = 0.4

b) Cost of equity, rs = Rf + beta x (Rm - Rf) = 6% + 1.25 x (11% - 6%) = 12.25%

Cost of debt, rd = 6%

WACC = wd x rd + ws x rs = 0.4 / 1.4 x 6% + 1 / 1.4 x 12.25% = 10.46%

c) Unlevered beta = Levered Beta / (1 + D/E) = 1.25 / (1 + 0.4) = 0.89

Cost of equity = Rf + beta x (Rm - Rf) = 6% + 0.89 x (11% - 6%) = 10.46%